A lightning bolt may carry a current of 1.00 × 10power5 A for short period of time. Find the magnetic field 50 m from the bolt. Assume that the bolt extends far above and below the point of observation.
1
Expert's answer
2013-08-16T11:21:01-0400
The formula for magnetic field created by the wire with currentI at distance r is
B=mu*I/(2*pi*r)
where mu = 4*pi*10^(-7) N*A^(2)
Hence we have
B = 2*10^(-7)*10^(5)/50 = 0.4*10^(-3) T
T means tesla, SI units for magnetic field